OpenStudy (shelby1290):

Factor completely using the most appropriate method. a) w^2-30-7w b)24x-x^3-10x^2 c)9-(2x+3)^2 d)x^2 - 1/4

OpenStudy (dayakar):

a)\[w^{2}+3w-10w-30\]

OpenStudy (dayakar):

w(w+3)-10(w+3)

OpenStudy (dayakar):

(w+3)(w-10)

OpenStudy (dayakar):

b)take -x common

OpenStudy (shelby1290):

for part A, how did you get 3 and -10? i thought you were supposed to find two numbers that multiply to -7 and add up to -30 3 X -10=-30 3 + (-10) =-7

OpenStudy (dayakar):

\[= -x(-24+x^{2}+10x)\]

OpenStudy (dayakar):

you are right

OpenStudy (dayakar):

\[=-x(x^{2}+10x-24)\]

OpenStudy (dayakar):

\[=-x(x^{2}-6x-4x+24)\]

OpenStudy (dayakar):

=-x[x(x-6)-4(x-6)

OpenStudy (dayakar):

=-x[(x-6)(x-4)]

OpenStudy (dayakar):

are u following

OpenStudy (shelby1290):

yes i am

OpenStudy (dayakar):

c)\[=3 ^{2}-\left( 2x+3 \right){2}\]

OpenStudy (shelby1290):

for part C, do you have to use the foil method?

OpenStudy (dayakar):

it in the form of \[x^{2}-y^{2}=(x+y)(x-y)\]

OpenStudy (dayakar):

[(3+(2x+3)][3-(2x+3)]

OpenStudy (dayakar):

(3+2x+3)(3-2x-3)

OpenStudy (dayakar):

(6+2x)(-2x)

OpenStudy (dayakar):

take 2 as common

OpenStudy (dayakar):

2(3+x)(-2x)

OpenStudy (dayakar):

=2(-2)x(x+3) =-4x(x+3)

OpenStudy (dayakar):

is it clear

OpenStudy (shelby1290):

so -4x(x+3) is the final answer? in my book, the answer is (2x+6)(-2x)

OpenStudy (dayakar):

both are same ,i simplified further

OpenStudy (dayakar):

ok

OpenStudy (dayakar):

d) is like above problem,can u try or may i help u

OpenStudy (shelby1290):

i'll try it and you can make sure that it's right

OpenStudy (dayakar):

good

OpenStudy (shelby1290):